OpenStudy (anonymous):

What is the approximate circumference of a circle with a diameter of feet? Use Pi = 3.14 A. 235.5 ft B. 471 ft C. 942 ft <------------------------ D. 2826 ft I spelled (Pi) not (pie) this time just for you @Unanimoose and @Bookworm14

OpenStudy (anonymous):

@Unanimoose @Bookworm14

OpenStudy (anonymous):

I think it is c

OpenStudy (anonymous):

I can help u don't worry

OpenStudy (anonymous):

thanks

OpenStudy (bookworm14):

how many feet is it?

OpenStudy (anonymous):

c is the right answer

OpenStudy (anonymous):

300

OpenStudy (anonymous):

sorry guys

OpenStudy (anonymous):

I got it I can help

OpenStudy (anonymous):

Circum = pi * Diam for 18 feet radius C = 3.14 * 18 * 2 = 113.04 feet for 500 km diameter C = 3.14 * 500 = 1570 km

OpenStudy (anonymous):



OpenStudy (anonymous):

up123 942 is right

OpenStudy (bookworm14):

so it would be 300 * 3.14 = 942, so yes you are correct!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

medal and fan me plz

OpenStudy (anonymous):

I will tag u again to give u that (I gave it to someone else first)